One number is 3 less than another. If 10 times the smaller number minus 4 times the larger numberis 24, find the two numbers.
Answer by stanbon(48558)

One number is 3 less than another. If 10 times the smaller number minus 4 times the larger numberis 24, find the two numbers.
------------
Let one of the numbers be "x" ; Then the smaller number is "x-3".
EQUATION:
10(x-3)-4x=24
10x-30-4x=24
6.=54
x=9 (the larger number)
x-3=6 (the smaller number)
-----------
cheers,
Stan H.